OpenStudy (anonymous):

changing y=x^2-8x+19 from standard form to vertex form

OpenStudy (anonymous):

Complete the square!!!

OpenStudy (anonymous):

y = (x^2 -8x +_______) +19 - __________ (-8/2)^2 =16 y = (x^2 -8x +___16____) +19 - ____16______ y = (x-4)(x-4) + 3 y = (x-4)^2 +3